# Rigidity Results for Compact Submanifolds with Pinched Ricci Curvature in Euclidean and Spherical Space Forms

### v2

@@ Abstract.
J. Q. Ge is partially supported by NSFC (No. 12171037) and the Fundamental Research Funds for the Central Universities.
Y. Zhou is partially supported by NSFC (No. 12171037, 12271040), China Postdoctoral Science Foundation (No. BX20230018) and National Key R&\&&D Program of China 2020YFA0712800.
@@ 1. Introduction
In 1979, Ejiri [6] obtained a celebrated rigidity theorem for $n(\geq 4)$-dimensional, simply connected, compact, orientable, minimal immersed submanifolds in the unit sphere with pinched Ricci curvature $\operatorname{Ric}\geq n-2$. In 2011, Xu-Tian [16] removed the assumption that the submanifold is simply connected. Later, Xu-Gu [14] generalized Ejiri’s rigidity theorem to $n(\geq 3)$-dimensional, compact, orientable submanifolds in a real space form $F^{n+m}(c)$ with $c+H^{2}>0$ under the assumption $\operatorname{Ric}\geq(n-2)\left(c+H^{2}\right)$ and parallel mean curvature vector field $\mathcal{H}$ ($H:=|\mathcal{H}|$). In the same paper, Xu-Gu raised a conjecture, indicating that the above generalization should also hold without the assumption of parallel mean curvature vector field. Recently, Dajczer-Vlachos [3] affirmed Xu-Gu’s conjecture for the case $n\geq 4$, expect for the part of diffeomorphic to spheres. The same philosophy was used in a series of rigidity results [1, 2, 4] for submanifolds with pinched Ricci curvature in different settings. For the case of spherical space forms [1, 2], the core techniques are different improvements of the following theorem:
@@
(((([12, Theorem 2])))) Let $M^{n}$ be an $n(\geq 4)$-dimensional compact submanifold of the unit sphere $\mathbf{S}^{n+m}$. Assume that the Ricci curvature satisfies
@@
Let $f:M^{n}\rightarrow F^{n+m}(c)$, $n\geq 5$, $c\geq 0$, be an isometric immersion of a compact, connected Riamannian manifold. If the Ricci curvature of $M$ satisfies
@@ 2. Preliminaries
Next, we will present several lemmas used in the proof of Theorem 1.2 and Corollary 1.3. The following Lemma 2.1 was first proved by Lawson-Simons [8] for the case $c>0$, and then by Xin [13] for the case $c=0$ when the inequality (##\##2.1) is strict on the whole submanifold. Elworthy-Rosenberg [7] observed that the result still holds by only requiring the inequality (##\##2.1) is strict at one point.
@@
(((([7, 8, 13])))). Let $f:M^{n}\rightarrow F^{n+m}(c)$, $c\geq 0$, be an isometric immersion of a compact manifold. Suppose that
(##\##) $\Theta_{q}:=\sum_{i=1}^{q}\sum_{j=q+1}^{n}\left(2|\operatorname{II}(e_{i},e_{j})|^{2}-\langle\operatorname{II}(e_{i},e_{i}),\operatorname{II}(e_{j},e_{j})\rangle\right)\leq q(n-q)c$
holds for any point $x\in M$ and any orthonormal basis $\{e_{1},\cdots,e_{n}\}$ of $T_{x}M$, where $q$ is an integer such that $11\leq q\leq n-$ 1. If there exists a point of $M$ such that the inequality ((((#italic-#\##2.1)))) is strict, then $H_{q}(M;\mathbb{Z})=H_{n-q}(M;\mathbb{Z})=0$.
@@
(((([3, Lemma 4])))) Let $f:M^{n}\rightarrow F^{n+m}(c)$, $n\geq 4$, $c\geq 0$, be an isometric immersion of a compact manifold. Suppose that the Ricci curvature of $M$ satisfies
@@
(((([10, Lemma 5])))) Let $f:M^{n}\rightarrow F^{n+m}(c)$ be a $1$-regular immersion. If for every point of $M$, there exists an orthonormal frame $\{E_{1},\cdots,E_{n}\}$ such that for each $1\leq i\leq n$, $\operatorname{II}(E_{i},E_{j})=0$ for any $j\neq i$ and $\operatorname{II}(E_{i},E_{i})=\sum_{j\neq i}a_{j}\operatorname{II}(E_{j},E_{j})$ for some real numbers $a_{j}$’s, then the first normal bundle $N_{1}$ is parallel.
@@
    The inequality ((((#italic-#\##2.1)))) is satisfied for $2\leq q\leq k$ at $x$ and is strict for $2\leq q<k$ if $c+H^{2}>0$ at $x$. Moreover, if the inequality ((((*2.5)))) is strict at $x$, then the inequality ((((#italic-#\##2.1)))) is strict for $q=k$ at $x$.
@@
    Suppose that the equality holds in ((((#italic-#\##2.1)))) with $q=k$ for an orthonormal basis $\{e_{1},\cdots,e_{n}\}$ of $T_{x}M$. Then for any orthonormal basis $\{\xi_{1},\cdots,\xi_{m}\}$ of $N_{x}M$, the corresponding matrix
@@
    where $\lambda_{\alpha}$ and $\mu_{\alpha}$ are principal curvatures with respect to $\xi_{\alpha}$ of multiplicities $k$ and $n-k$, respectively. Furthermore, if $c+H^{2}>0$ at $x$, then $\eta_{1}=\sum_{\alpha=1}^{m}\lambda_{\alpha}\xi_{\alpha}$ and $\eta_{2}=\sum_{\alpha=1}^{m}\mu_{\alpha}\xi_{\alpha}$ are distinct principal normals of $f$ at $x$ such that $E_{\eta_{1}}(x)=\operatorname{Span}\{e_{1},\cdots,e_{k}\}$ and $E_{\eta_{2}}(x)=\operatorname{Span}\{e_{k+1},\cdots,e_{n}\}$. In addition,
@@
for any $2\leq q\leq k$. Note that the last inequality in (2.11) is strict for $2\leq q<k$ if $c+H^{2}>0$ at $x$. Hence the inequality (##\##2.1) is satisfied for $2\leq q\leq k$ at $x$ and is strict for $2\leq q<k$ if $c+H^{2}>0$ at $x$. If in addition the inequality (*2.5) is strict at $x$, then the second inequality in (2.11) is strict, and thus the inequality (##\##2.1) is strict for $q=k$ at $x$.
Next, we prove $(2)$. Assume that the equality holds in (##\##2.1) with $q=k$ for an orthonormal basis $\{e_{1},\cdots,e_{n}\}$ of $T_{x}M$. Then all inequalities in (2.9-2.11) become equalities. The first inequality in (2.10) implies that for each $1\leq\alpha\leq m$, there exist $\lambda_{\alpha},\mu_{\alpha}\in\mathbb{R}$ such that
@@ 3. Proof of Theorems
Next, it follows immediately from Lemma 2.5 (1) that the inequality (##\##2.1) is satisfied for $2\leq q\leq k$ at every point of $M$ and strict for $2\leq p<k$ at points with $c+H^{2}>0$. Then Lemma 2.1 implies that
@@
Case (2): $H_{k}(M;\mathbb{Z})=H_{n-k}(M;\mathbb{Z})=0$ does not hold. It follows from Lemma 2.1 that for any point $x\in M$, there exists an orthonormal basis $\{e_{1},\cdots,e_{n}\}$ of $T_{x}M$ such that the equality holds in (##\##2.1) for $q=k$. Then by Lemma 2.5 (2), for any smooth orthonormal local frame $\left(\xi_{1},\cdots,\xi_{m}\right)$ for $NM$ over an open subset $U$ of $M$, the corresponding matrix
@@
where $\lambda_{\alpha}$ and $\mu_{\alpha}$ are principal curvatures with respect to $\xi_{\alpha}$ of multiplicities $k$ and $n-k$, respectively. It follows that the normal bundle $NM$ is flat. In addition, by Lemma 2.5 (2), we have
@@
Case (2-a): $\operatorname{rank}N_{1}=1$. In this case, it follows from (3.4-3.7) that $\lambda_{1}\mu_{1}\neq 0$ as $H>0$ and $\lambda_{\alpha}=\mu_{\alpha}=0$ for each $2\leq\alpha\leq m$. Then $\eta_{1}=\lambda_{1}\xi_{1}$ and $\eta_{2}=\mu_{1}\xi_{1}$ are smooth normal vector fields and hence Dupin principal normal vector fields with multiplicities $k$ and $n-k$, respectively. Therefore, $E_{\eta_{1}}$ and $E_{\eta_{2}}$ are smooth distributions such that $TM=E_{\eta_{1}}\oplus E_{\eta_{2}}$. For every point of $M$, we can choose a smooth orthonormal local frame $\left(E_{1},\cdots,E_{n}\right)$ such that $E_{\eta_{1}}=\operatorname{Span}\{E_{1},\cdots,E_{k}\}$ and $E_{\eta_{2}}=\operatorname{Span}\{E_{k+1},\cdots,E_{n}\}$. It follows that
@@
Case (2-b): $\operatorname{rank}N_{1}=2$. In this case, we can assume additionally that $N_{1}\big{|}_{U}=\operatorname{Span}\{\xi_{1},\xi_{2}\}$. Then we have $\eta_{1}=\lambda_{1}\xi_{1}+\lambda_{2}\xi_{2}$ and $\eta_{2}=\mu_{1}\xi_{1}+\mu_{2}\xi_{2}$. It follows from (3.5-3.9) that $\lambda_{2},\mu_{2}$ are non-zero constants. Then $\eta_{1}$ and $\eta_{2}$ are smooth normal vector fields and hence Dupin principal normal vector fields with multiplicities $k$ and $n-k$, respectively. Similarly, $E_{\eta_{1}}$ and $E_{\eta_{2}}$ are smooth distributions such that $TM=E_{\eta_{1}}\oplus E_{\eta_{2}}$. For every point of $M$, we can choose a smooth orthonormal local frame $\left(E_{1},\cdots,E_{n}\right)$ such that $E_{\eta_{1}}=\operatorname{Span}\{E_{1},\cdots,E_{k}\}$ and $E_{\eta_{2}}=\operatorname{Span}\{E_{k+1},\cdots,E_{n}\}$. Note that (3.2) implies that
@@
It follows from Lemma 2.4 that $N_{1}$ is parallel, i.e., $\nabla_{X}^{\bot}\xi\subset N_{1}$ for any $X\in TM$ and any smooth section $\xi$ of $N_{1}$. By $\langle\xi_{1},\xi_{1}\rangle=\langle\xi_{2},\xi_{2}\rangle=1$, we obtain
@@
such that $f=u\circ g$ and $\alpha(n,k,H,c)=(n-l)\overline{c}$ for an positive integer $l$, where $r_{i}^{2}=\frac{n_{i}-1}{\alpha(n,k,H,c)}$, $2\leq n_{i}\leq n$ and $\sum_{i=1}^{l}n_{i}=n$. Since $u$ is totally umbilical, it is easy to verify that $H^{2}=H_{g}^{2}+H_{u}^{2}$ and $\overline{c}=c+H_{u}^{2}$, where $H_{g}$ and $H_{u}$ denote the mean curvature of $g$ and $u$, respectively. It follows that $\alpha(n,k,H,c)=(n-l)\left(c+H_{u}^{2}\right)\leq(n-l)\left(c+H^{2}\right)$. Then by Lemma 2.3, we have $n-3<n-l$, i.e., $l=1$ or $2$. Recall that $H_{k}(M;\mathbb{Z})=H_{n-k}(M;\mathbb{Z})=0$ does not hold in Case (2). Hence, we must have $l=2$ and $M$ is isometric to the Einstein hypersurface $S^{k}\left(\sqrt{\frac{k-1}{n-2}}\,r\right)\times S^{n-k}\left(\sqrt{\frac{n-k-1}{n-2}}\,r\right)$ in the totally umbilical sphere $S^{n+1}\left(r\right)\subset F^{n+m}(c)$, where $r=\sqrt{\frac{n-2}{\alpha(n,k,H,c)}}$. Finally, one can verify by a direct calculation that the Ricci curvature and the mean curvature of this Einstein hypersurface are exactly
@@ Acknowledgments.
Data Availibility.
This manuscript has no associated data.
Conflict of interest.
On behalf of all authors, the corresponding author states that there is no conflict of interest.

### v3

@@ Abstract.
J. Q. Ge is partially supported by NSFC (No. 12171037, 12571049) and the Fundamental Research Funds for the Central Universities.
Y. Zhou is partially supported by NSFC (No. 12171037, 12271040), China Postdoctoral Science Foundation (No. BX20230018) and National Key R&\&D Program of China 2020YFA0712800.
@@ 1. Introduction
In 1979, Ejiri [6] obtained a celebrated rigidity theorem for $n(\geq 4)$-dimensional, simply connected, compact, orientable, minimal immersed submanifolds in the unit sphere with pinched Ricci curvature $\operatorname{Ric}\geq n-2$. In 2011, Xu-Tian [16] removed the assumption that the submanifold is simply connected. Later, Xu-Gu [14] generalized Ejiri’s rigidity theorem to $n(\geq 3)$-dimensional, compact, orientable submanifolds in a real space form $F^{n+m}(c)$ with $c+H^{2}>0$ under the assumption $\operatorname{Ric}\geq(n-2)\left(c+H^{2}\right)$ and parallel mean curvature vector field $\mathcal{H}$ ($H:=|\mathcal{H}|$). In the same paper, Xu-Gu raised a conjecture, indicating that the above generalization should also hold without the assumption of parallel mean curvature vector field. Recently, Dajczer-Vlachos [3] affirmed Xu-Gu’s conjecture for the case $n\geq 4$, except for the part of diffeomorphic to spheres. The same philosophy was used in a series of rigidity results [1, 2, 4] for submanifolds with pinched Ricci curvature in different settings. For the case of spherical space forms [1, 2], the core techniques are different improvements of the following theorem:
@@
(([12, Theorem 2])) Let $M^{n}$ be an $n(\geq 4)$-dimensional compact submanifold of the unit sphere $\mathbf{S}^{n+m}$. Assume that the Ricci curvature satisfies
@@ 2. Preliminaries
Next, we will present several lemmas used in the proof of Theorem 1.2 and Corollary 1.3. The following Lemma 2.1 was first proved by Lawson-Simons [8] for the case $c>0$, and then by Xin [13] for the case $c=0$ when the inequality (#\#2.1) is strict on the whole submanifold. Elworthy-Rosenberg [7] observed that the result still holds by only requiring the inequality (#\#2.1) is strict at one point.
@@
(([7, 8, 13])). Let $f:M^{n}\rightarrow F^{n+m}(c)$, $c\geq 0$, be an isometric immersion of a compact manifold. Suppose that
(#\#) $\Theta_{q}:=\sum_{i=1}^{q}\sum_{j=q+1}^{n}\left(2|\operatorname{II}(e_{i},e_{j})|^{2}-\langle\operatorname{II}(e_{i},e_{i}),\operatorname{II}(e_{j},e_{j})\rangle\right)\leq q(n-q)c$
holds for any point $x\in M$ and any orthonormal basis $\{e_{1},\cdots,e_{n}\}$ of $T_{x}M$, where $q$ is an integer such that $1\leq q\leq n-1$. If there exists a point of $M$ such that the inequality ((#\#2.1)) is strict, then $H_{q}(M;\mathbb{Z})=H_{n-q}(M;\mathbb{Z})=0$.
@@
(([3, Lemma 4])) Let $f:M^{n}\rightarrow F^{n+m}(c)$, $n\geq 4$, $c\geq 0$, be an isometric immersion of a compact manifold. Suppose that the Ricci curvature of $M$ satisfies
@@
(([10, Lemma 5])) Let $f:M^{n}\rightarrow F^{n+m}(c)$ be a $1$-regular isometric immersion. If for every point of $M$, there exists an orthonormal frame $\{E_{1},\cdots,E_{n}\}$ such that for each $1\leq i\leq n$, $\operatorname{II}(E_{i},E_{j})=0$ for any $j\neq i$ and $\operatorname{II}(E_{i},E_{i})=\sum_{j\neq i}a_{j}\operatorname{II}(E_{j},E_{j})$ for some real numbers $a_{j}$’s, then the first normal bundle $N_{1}$ is parallel.
@@
    The inequality ((#\#2.1)) is satisfied for $2\leq q\leq k$ at $x$ and is strict for $2\leq q<k$ if $c+H^{2}>0$ at $x$. Moreover, if the inequality ((*2.5)) is strict at $x$, then the inequality ((#\#2.1)) is strict for $q=k$ at $x$.
@@
    Suppose that the equality holds in ((#\#2.1)) with $q=k$ for an orthonormal basis $\{e_{1},\cdots,e_{n}\}$ of $T_{x}M$. Then for any orthonormal basis $\{\xi_{1},\cdots,\xi_{m}\}$ of $N_{x}M$, the corresponding matrix
@@
    where $\lambda_{\alpha}$ and $\mu_{\alpha}$ are principal curvatures with respect to $\xi_{\alpha}$ of multiplicities $k$ and nkn-k, respectively. Furthermore, if $c+H^{2}>0$ at $x$, then $\eta_{1}=\sum_{\alpha=1}^{m}\lambda_{\alpha}\xi_{\alpha}$ and $\eta_{2}=\sum_{\alpha=1}^{m}\mu_{\alpha}\xi_{\alpha}$ are distinct principal normals of $f$ at $x$ such that $E_{\eta_{1}}(x)=\operatorname{Span}\{e_{1},\cdots,e_{k}\}$ and $E_{\eta_{2}}(x)=\operatorname{Span}\{e_{k+1},\cdots,e_{n}\}$. In addition,
@@
for any $2\leq q\leq k$. Note that the last inequality in (2.11) is strict for $2\leq q<k$ if $c+H^{2}>0$ at $x$. Hence the inequality (#\#2.1) is satisfied for $2\leq q\leq k$ at $x$ and is strict for $2\leq q<k$ if $c+H^{2}>0$ at $x$. If in addition the inequality (*2.5) is strict at $x$, then the second inequality in (2.11) is strict, and thus the inequality (#\#2.1) is strict for $q=k$ at $x$.
Next, we prove $(2)$. Assume that the equality holds in (#\#2.1) with $q=k$ for an orthonormal basis $\{e_{1},\cdots,e_{n}\}$ of $T_{x}M$. Then all inequalities in (2.9-2.11) become equalities. The first inequality in (2.10) implies that for each $1\leq\alpha\leq m$, there exist $\lambda_{\alpha},\mu_{\alpha}\in\mathbb{R}$ such that
@@ 3. Proof of Theorems
Next, it follows immediately from Lemma 2.5 (1) that the inequality (#\#2.1) is satisfied for $2\leq q\leq k$ at every point of $M$ and strict for $2\leq p<k$ at points with $c+H^{2}>0$. Then Lemma 2.1 implies that
@@
Case (2): $H_{k}(M;\mathbb{Z})=H_{n-k}(M;\mathbb{Z})=0$ does not hold. It follows from Lemma 2.1 that for any point $x\in M$, there exists an orthonormal basis $\{e_{1},\cdots,e_{n}\}$ of $T_{x}M$ such that the equality holds in (#\#2.1) for $q=k$. Then by Lemma 2.5 (2), for any smooth orthonormal local frame $\left(\xi_{1},\cdots,\xi_{m}\right)$ for $NM$ over an open subset $U$ of $M$, the corresponding matrix
@@
where $\lambda_{\alpha}$ and $\mu_{\alpha}$ are principal curvatures with respect to $\xi_{\alpha}$ of multiplicities $k$ and nkn-k, respectively. It follows that the normal bundle $NM$ is flat. In addition, by Lemma 2.5 (2), we have
@@
Case (2-a): $\operatorname{rank}N_{1}=1$. In this case, it follows from (3.4-3.7) that $\lambda_{1}\mu_{1}\neq 0$ as $H>0$ and $\lambda_{\alpha}=\mu_{\alpha}=0$ for each $2\leq\alpha\leq m$. Then $\eta_{1}=\lambda_{1}\xi_{1}$ and $\eta_{2}=\mu_{1}\xi_{1}$ are smooth normal vector fields and hence Dupin principal normal vector fields with multiplicities $k$ and nkn-k, respectively. Therefore, $E_{\eta_{1}}$ and $E_{\eta_{2}}$ are smooth distributions such that $TM=E_{\eta_{1}}\oplus E_{\eta_{2}}$. For every point of $M$, we can choose a smooth orthonormal local frame $\left(E_{1},\cdots,E_{n}\right)$ such that $E_{\eta_{1}}=\operatorname{Span}\{E_{1},\cdots,E_{k}\}$ and $E_{\eta_{2}}=\operatorname{Span}\{E_{k+1},\cdots,E_{n}\}$. It follows that
@@
Case (2-b): $\operatorname{rank}N_{1}=2$. In this case, we can assume additionally that $N_{1}\big|_{U}=\operatorname{Span}\{\xi_{1},\xi_{2}\}$. Then we have $\eta_{1}=\lambda_{1}\xi_{1}+\lambda_{2}\xi_{2}$ and $\eta_{2}=\mu_{1}\xi_{1}+\mu_{2}\xi_{2}$. It follows from (3.5-3.9) that $\lambda_{2},\mu_{2}$ are non-zero constants. Then $\eta_{1}$ and $\eta_{2}$ are smooth normal vector fields and hence Dupin principal normal vector fields with multiplicities $k$ and nkn-k, respectively. Similarly, $E_{\eta_{1}}$ and $E_{\eta_{2}}$ are smooth distributions such that $TM=E_{\eta_{1}}\oplus E_{\eta_{2}}$. For every point of $M$, we can choose a smooth orthonormal local frame $\left(E_{1},\cdots,E_{n}\right)$ such that $E_{\eta_{1}}=\operatorname{Span}\{E_{1},\cdots,E_{k}\}$ and $E_{\eta_{2}}=\operatorname{Span}\{E_{k+1},\cdots,E_{n}\}$. Note that (3.2) implies that
@@
It follows from Lemma 2.4 that $N_{1}$ is parallel, i.e., $\nabla_{X}^{\bot}\xi\in\Gamma(N_{1})$ for any $X\in TM$ and any smooth section $\xi\in\Gamma(N_{1})$. By $\langle\xi_{1},\xi_{1}\rangle=\langle\xi_{2},\xi_{2}\rangle=1$, we obtain
@@
such that $f=u\circ g$ and $\alpha(n,k,H,c)=(n-l)\overline{c}$ for an positive integer $l$, where $r_{i}^{2}=\frac{n_{i}-1}{\alpha(n,k,H,c)}$, $2\leq n_{i}\leq n$ and $\sum_{i=1}^{l}n_{i}=n$. Since $u$ is totally umbilical, it is easy to verify that $H^{2}=H_{g}^{2}+H_{u}^{2}$ and $\overline{c}=c+H_{u}^{2}$, where $H_{g}$ and $H_{u}$ denote the mean curvature of $g$ and $u$, respectively. It follows that $\alpha(n,k,H,c)=(n-l)\left(c+H_{u}^{2}\right)\leq(n-l)\left(c+H^{2}\right)$. Then by Lemma 2.3, we have n3<nln-3<n-l, i.e., $l=1$ or $2$. Recall that $H_{k}(M;\mathbb{Z})=H_{n-k}(M;\mathbb{Z})=0$ does not hold in Case (2). Hence, we must have $l=2$ and $M$ is isometric to the Einstein hypersurface $S^{k}\left(\sqrt{\frac{k-1}{n-2}}\,r\right)\times S^{n-k}\left(\sqrt{\frac{n-k-1}{n-2}}\,r\right)$ in the totally umbilical sphere $S^{n+1}\left(r\right)\subset F^{n+m}(c)$, where $r=\sqrt{\frac{n-2}{\alpha(n,k,H,c)}}$. Finally, one can verify by a direct calculation that the Ricci curvature and the mean curvature of this Einstein hypersurface are exactly